# Perturbation-Robust Predictive Modeling of Social Effects by Network Subspace Generalized Linear Models

### v4

@@ Abstract
Network-linked data, where multivariate observations are interconnected by a network, are becoming increasingly prevalent in fields such as sociology and biology. These data often exhibit inherent noise and complex relational structures, complicating conventional modeling and statistical inference. Motivated by empirical challenges in analyzing such data sets, this paper introduces a family of network subspace generalized linear models designed for analyzing noisy, network-linked data. We propose a model inference method based on subspace-constrained maximum likelihood, which emphasizes flexibility in capturing network effects and provides a robust inference framework against network perturbations. We establish the asymptotic distributions of the estimators under network perturbations, demonstrating the method’s accuracy through extensive simulations involving random network models and deep-learning-based embedding algorithms. The proposed methodology is applied to a comprehensive analysis of a large-scale study on school conflicts, where it identifies significant social effects, offering meaningful and interpretable insights into student behaviors.
@@ 1 Introduction
Network data analysis has become increasingly popular due to its wide-ranging applications in the social sciences (Holme, 2015; Van den Bos et al., 2018), biological sciences (Özgür et al., 2008; Zeng et al., 2018), and engineering (Le Gat, 2014; Cuadra et al., 2015). A notable category of social network data concerns network-linked objects, where the interactions or relationships among individuals are depicted through network structures, and each individual typically has associated response variables and covariates. Such structures are frequently encountered in studies examining social influences on human behavior (Michell and West, 1996; Michell, 2000; Harris, 2009; Paluck et al., 2016). In this paper, we focus on analyzing student behavior in the context of school conflicts, using the data from Paluck et al. (2016). Despite the development of numerous statistical models to analyze network-linked data in recent years (Zhang et al., 2016; Li et al., 2019; Su et al., 2019; Zhang et al., 2020; Sit and Ying, 2021; Mao et al., 2021; Mukherjee et al., 2021; Le and Li, 2022; Hayes et al., 2022; Fang et al., 2023; He et al., 2023; Lunde et al., 2023; Zhu et al., 2017; Wu and Leng, 2023; Armillotta and Fokianos, 2023; Chang and Paul, 2024), the noisy nature of the network structures in this study necessitates non-trivial generalizations of the existing literature to effectively analyze the school conflict data. This challenge motivates the development of our new model. In the following sections, we introduce the school conflict study and review the current literature on predictive modeling for network-linked data.
1.1 Social effect analysis in the school conflicts study
@@
A prospective study by Paluck et al. (2016) investigated the effects of randomized anti-conflict interventions on social norms across 56 high schools in New Jersey. In the treatment schools, students were randomly selected to participate in educational workshops on school conflicts. Data collection included official records from school administrations and student questionnaires, where students provided personal information, opinions on conflict-related events, and a list of their closest friends at both the beginning and end of the school year, allowing for the mapping of social networks within each school. In 25 of these schools, which were randomly selected from the 56, the experimenters introduced educational workshops aimed at a small group of students to reduce school conflicts. The field experiment sought to demonstrate that introducing educational interventions to students could help mitigate conflicts within schools. The anti-conflict impact was measured through the allocation of orange wristbands, which rewarded students for friendly or conflict-mitigating actions.
@@
In this study, a key object of interest is the social influence, which could play a significant role in disseminating the effects of the intervention throughout the entire school. To facilitate the analysis of social influence, the experimenters recorded friendship relations in terms of “how much time two students spend together.” In addition to social influence, the study aims to understand the impact of various background covariates, such as gender, race, and family conditions. While the original study by Paluck et al. (2016) utilized social relations to infer social effects, recent work by Le and Li (2022) highlighted the importance of accounting for noisy observations in friendship relations to ensure valid inference. Specifically, two waves of surveys were administered within the same school year to capture social relations. However, the overlap between the two waves was limited. Figure 1 displays the edge overlap proportions across the 25 schools, showing that, in most schools, only about 50% of the edges overlapped between the two periods. Such high levels of noise in the observations can jeopardize the validity of statistical inference if the network structure errors are not adequately addressed in the analysis.
@@
Figure 1: Left panel: the proportion of overlapping edges in two waves in the study, across 25 schools. Right panel: the networks in two waves in one example school.
More generally, noisy observations of network structures are frequently encountered in other empirical studies, particularly in the social sciences (Onnela et al., 2007; Yu et al., 2008; Harris, 2009). Furthermore, popular graph-embedding methods in machine learning (Perozzi et al., 2014; Grover and Leskovec, 2016; Rozemberczki and Sarkar, 2018), which are commonly used to manage network data in modeling tasks, may introduce additional perturbations due to their inherent randomness. These challenges underscore the need for a general predictive modeling strategy that can account for network perturbations while ensuring validity in inference. Addressing this issue is the central focus of our model development in this paper.
@@ 1.2 Predictive models for network-linked data
We focus on predictive modeling for a node-level response variable. Among the available works in literature, relevant predictive models can be categorized into three main classes based on their design. The first class treats the network as a generalized spatial relationship and employs a graph-based autoregressive model to capture dependencies (Zhu et al., 2017; Armillotta and Fokianos, 2023; Wu and Leng, 2023; Chang and Paul, 2024). The second class of methods incorporates network information through a distance-based dependence structure, assuming that responses are independent if the distance between them exceeds a certain threshold (Su et al., 2019; Sit and Ying, 2021; Mukherjee et al., 2021). Both of these classes rely on parametric forms of network effects and assume that the observed networks are accurate. While these methods provide informative model inference if the assumed network effect is appropriate, they may lead to misleading conclusions when the assumed parametric form is violated. Additionally, they tend to be vulnerable to errors in the observed network structure, which is a key issue in our motivating application.
The third class of methods employs nonparametric components to model network effects. For instance, Li et al. (2019) introduces the regression with network cohesion (RNC) approach, which includes an individual node effects component along with a network smoothing penalty. This method has proven to be flexible for modeling network-linked responses and is applicable to various settings, such as generalized linear models. However, this approach lacks a formal statistical inference framework. A more recent model in this category is the subspace linear regression proposed by Le and Li (2022). Instead of assuming a smooth network effect, this model posits a latent subspace for the network effect. It offers a valid inference framework and demonstrates robustness against network perturbations. However, the model fitting relies on a sequence of geometric projections, which are valid only for linear regression. This restriction can be limiting in practice, as categorical and discrete responses are common in social science applications, such as our motivating example of the school conflict study. In a separately line, Hayes et al. (2022) introduced a model in a similar flavor to to analyze network-mediated effect in causal problems, in which the network effect is parameterized by linear combination of latent vectors. This method can handle other types of response variables, but the latent vectors are assumed to follow the random dot product graph model (Athreya et al., 2018).
@@ 2.1 Notations
Throughout the paper, we use $c,C>0$ to denote absolute constants, the values of which may change from line to line. For two sequences of positive scalars $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$, we write $a_{n}=o(b_{n})$ and $a_{n}=O(b_{n})$ if $a_{n}/b_{n}$ converges to zero and $a_{n}/b_{n}$ is bounded, respectively. Similarly, for a sequence of random variables $\{X_{n}\}_{n=1}^{\infty}$, we write $X_{n}=o_{p}(b_{n})$ and $X_{n}=O_{p}(b_{n})$ if $X_{n}/b_{n}$ converges to zero and is bounded in probability, respectively. We use $I_{n}\in\mathbb{R}^{n\times n}$ to denote the identity matrix of size $n$. For a matrix $A=(A_{ij})\in\mathbb{R}^{n\times n}$, $\mathrm{tr}(A)=\sum_{i=1}^{n}A_{ii}$ is the trace, while $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ are the minimum and maximum eigenvalues of $A$, respectively, when $A$ is symmetric. For a vector $u$, $\|u\|$ is the Euclidean norm. For a matrix $W=(W_{ij})\in\mathbb{R}^{m\times n}$ with $1\leq n\leq m$ and the singular decomposition $W=\sum_{i=1}^{n}\sigma_{i}u_{i}v_{i}^{\top}$, $\|W\|=\max_{1\leq i\leq n}\sigma_{i}$, $\|W\|_{F}=(\sum_{i=1}^{n}\sigma^{2}_{i})^{1/2}$ and $\|W\|_{\infty}=\max_{1\leq i\leq n}\sum_{j=1}^{n}|W_{ij}|$ represent the spectral norm, the Frobenius norm, and the infinity norm of $A$, respectively. In addition, $W_{i}$ is the $i$-th column of $W$, and $W_{u:v}$ is the sub-matrix of $W$ with column vectors $W_{i}$ for $u\leq i\leq v$. We further use $W_{i,u:v}$ to denote the $i$-th row of $W_{u:v}$.
@@ 2.2 Model
We assume there exists a true unobserved relational matrix $P\in\mathbb{R}^{n\times n}$, where $P_{ij}$ describes the strength of the relationship between the nodes $i$ and $j$. Let $\hat{P}=(\hat{P}_{ij})\in\mathbb{R}^{n\times n}$ be an approximate relational matrix, which can be viewed as a noisy version of $P$ that is computable from observed relations between observations. An example of this relational model appears in random network modeling literature assumes that the entries of $\hat{P}$ are the observed adjacency connections between nodes, generated as independent Bernoulli random variables with $P=\mathbb{E}[\hat{P}]$, or some improved estimators based on certain statistical estimation methods (Li and Le, 2023); another example discussed in detail in Section 4.2 involves stochastic embedding algorithms for which $\hat{P}$ is the similarity between the random embedding output. Intuitively, we expect that $\hat{P}$ does not significantly deviate from $P$.
@@
Conditioning on $X$ and $P$, we assume that $y_{1},...,y_{n}$ are independent random variables drawn from a generalized linear model (GLM). Followng McCullagh (2019), the probability density or probability mass function of $y_{i}$ can be expressed in the following form:
@@
Here, $a$, $b$, and $d$ are specific functions depending on the distribution of $y_{i}$. For example, when $a(\phi)=1$, $b(\psi_{i})=\log(1+e^{\psi_{i}})$, and $d(y,\phi)=0$, (1) leads to a logistic regression; when $a(\phi)=1$, $b(\psi_{i})=e^{\psi_{i}}$, and $d(y,\phi)=-\log(y!)$, a Poisson regression is obtained. In addition, $\phi$ is a known dispersion parameter, and $\psi$ is a link function. We write
@@
through a link function, where $S_{K}(P)$ is the linear subspace spanned by the $K$ leading eigenvectors of $P$. The assumption that $\alpha$ belongs to $S_{K}(P)$ is natural and supported by existing evidence that leading eigenvectors of the relational matrix typically capture crucial network information (Özgür et al., 2008; Zeng et al., 2018; Van den Bos et al., 2018; Lee, 2019). In particular, building on the modeling approach outlined in Le and Li (2022), we assume that $\mu$ is contained in the Minkowski sum of $\operatorname{col}(X)$ and $S_{K}(P)$ through the link function $h^{-1}$, which is assumed to be smooth and increasing:
@@ 2.3 Model fitting by the subspace-constrained maximum likelihood
We now describe the model fitting procedure for the network subspace generalized linear model. For ease of presentation, let us first outline this procedure assuming we observe $S_{K}(P)$. At a high level, we want to use the restricted maximum likelihood estimator (MLE) under the subspace constraint under Definition 2.1. Therefore, the estimation is done by solving the following optimization problem:
@@
$\sigma_{1}=\sigma_{2}=\cdots=\sigma_{r}=1>\sigma_{r+1}\geq\cdots\geq\sigma_{r+s}>0=\sigma_{r+s+1}=\cdots=0.$ (6)
To calculate a basis for the intersection subspace $\mathcal{R}$, let us denote
@@
Viewing $g_{i}$ as a new covariate vector for the $i$-th observation turns the model of Definition 2.1 into a typical generalized linear model with parameter $\gamma^{*}$ (if we do know $g_{i}$’s). Using the first-order stationary condition and setting the gradient of the likelihood function to zero lead to the following estimating equation:
@@
where $h^{\prime}(\cdot)$ is the derivative of the link function and $v(g_{i}^{\top}\gamma)$ is the variance of $y_{i}$. Taking the partial derivative of $-S(\gamma)$, we obtain the oracle information matrix
@@
Sample version estimators. In practice, instead of observing the relational matrix $P$ directly, we only have access to a noisy version $\hat{P}$ of $P$. We replace $P$ with $\hat{P}$ everywhere in the above procedure. In particular, let $\breve{W}\in\mathbb{R}^{n\times K}$ be the matrix whose columns are eigenvectors of $\hat{P}$ that span the subspace $S_{K}(\hat{P})$. The singular value decomposition of $\bar{Z}^{\top}\breve{W}$ takes the form
$\bar{Z}^{\top}\breve{W}=\tilde{U}\tilde{\Sigma}\tilde{V}^{T}.$
@@
We always assume that $r$, the dimension of $\mathcal{R}$, is known, as a consistent estimate of $r$ is already provided by Le and Li (2022). With this assumption, we estimate $\mathcal{R}$, $\mathcal{C}$, and $\mathcal{N}$ by
^$\hat{\mathcal{R}}=\operatorname{col}(\tilde{Z}_{1:r}),\quad\hat{\mathcal{C}}=\operatorname{col}(\tilde{Z}_{(r+1):p}),\quad\hat{\mathcal{N}}=\operatorname{col}(\tilde{W}_{(r+1):K}).$ (15)
@@
1 Calculate the orthonormal basis of $\operatorname{col}(X)$ and form matrix $Z\in\mathbb{R}^{n\times p}$ in (7); calculate $K$ eigenvectors of $\hat{P}$ and form $\breve{W}\in\mathbb{R}^{n\times K}$.
2Calculate the singular value decomposition $Z^{T}\breve{W}=\tilde{U}\tilde{\Sigma}\tilde{V}^{T}$, and form $\tilde{Z}=\sqrt{n}Z\hat{U},\tilde{W}=\sqrt{n}{W}\hat{V}.$
Find the root $\hat{\gamma}$ of the generalized estimating equation $\tilde{S}(\gamma)=0$ using the iteratively reweighted least squares method, and obtain $\hat{\theta},\hat{\beta},\hat{\alpha}$ by replacing $\gamma$ with $\hat{\gamma}$ in (9), (10), and (11), respecitvely.
Algorithm 1 Subspace Maximum Likelihood Estimation Algorithm

### v6

@@ Abstract
Network-linked data, in which multivariate observations are interconnected by a network, are becoming increasingly prevalent in fields such as sociology and biology. These data often exhibit inherent noise and complex relational structures, complicating conventional modeling and statistical inference. Motivated by empirical challenges in analyzing such datasets, this paper introduces a family of network subspace generalized linear models designed for analyzing noisy, network-linked data. We propose a model inference method based on subspace-constrained maximum likelihood that emphasizes flexibility in capturing network effects and provides an inference framework that is robust under network perturbations. We establish the asymptotic distributions of the estimators under network perturbations, demonstrating the method’s accuracy through extensive simulations involving random network models and deep-learning-based embedding algorithms. The proposed methodology is applied to a comprehensive analysis of a large-scale study on school conflicts, where it identifies significant social effects, offering meaningful and interpretable insights into student behavior.
@@ 1 Introduction
Network data analysis has become increasingly popular due to its wide-ranging applications in the social sciences (Holme, 2015; Van den Bos et al., 2018), biological sciences (Özgür et al., 2008; Zeng et al., 2018), and engineering (Le Gat, 2014; Cuadra et al., 2015). A notable category of social network data concerns network-linked objects, in which the interactions or relationships among individuals are depicted through network structures, and each individual typically has associated response variables and covariates. Such structures are frequently encountered in studies examining social influences on human behavior (Michell and West, 1996; Michell, 2000; Harris, 2009; Paluck et al., 2016). In this paper, we focus on analyzing student behavior in the context of school conflicts, using data from Paluck et al. (2016). Despite the development of numerous statistical models to analyze network-linked data in recent years (Zhang et al., 2016; Li et al., 2019; Su et al., 2019; Zhang et al., 2020; Sit and Ying, 2021; Mao et al., 2021; Mukherjee et al., 2021; Le and Li, 2022; Hayes et al., 2022; Fang et al., 2023; He et al., 2023; Lunde et al., 2023; Zhu et al., 2017; Wu and Leng, 2023; Armillotta and Fokianos, 2023; Chang and Paul, 2024), the noisy nature of the network structures in this study necessitates non-trivial generalizations of the methods in the existing literature to effectively analyze the school conflict data. This challenge motivates the development of our new model. In the following sections, we introduce the school conflict study and review the current literature on predictive modeling for network-linked data.
1.1 Social effect analysis in the school conflict study
@@
A prospective study by Paluck et al. (2016) investigated the effects of randomized anti-conflict interventions on social norms across 56 high schools in New Jersey. Data collection included official records from school administrations and student questionnaires, where students provided personal information, opinions on conflict-related events, and a list of their closest friends at both the beginning and end of the school year, allowing for the mapping of social networks within each school. In 25 of these schools, which were randomly selected from the 56, the experimenters introduced educational workshops aimed at a small group of students to reduce school conflicts. The field experiment sought to demonstrate that introducing educational interventions to students could help mitigate conflicts within schools. The anti-conflict impact was measured through the distribution of orange wristbands, which rewarded students for friendly or conflict-mitigating actions.
@@
In this study, a key quantity of interest is the social influence, which could play a significant role in disseminating the effects of the intervention throughout the entire school. To facilitate the analysis of social influence, the experimenters recorded friendship relations in terms of “how much time two students spent together.” In addition to social influence, the study aims to understand the impact of various background covariates, such as gender, race, and family conditions. While the original study by Paluck et al. (2016) utilized social relations to infer social effects, recent work by Le and Li (2022) highlighted the importance of accounting for noisy observations in friendship relations to ensure valid inference. Specifically, two waves of surveys were administered within the same school year to capture social relations. However, the overlap between the two waves was limited. Figure 1 displays the edge overlap proportions across the 25 schools, showing that, in most schools, only about 50% of the edges overlapped between the two periods. Such high levels of noise in the observations can jeopardize the validity of statistical inference if the network structure errors are not adequately addressed in the analysis.
@@
Figure 1: Left panel: the proportion of overlapping edges in two waves in the study, across 25 schools. Right panel: the networks in two waves for one example school.
More generally, noisy observations of network structures are frequently encountered in other empirical studies, particularly in the social sciences (Onnela et al., 2007; Yu et al., 2008; Harris, 2009). Furthermore, popular graph-embedding methods in machine learning (Perozzi et al., 2014; Grover and Leskovec, 2016; Rozemberczki and Sarkar, 2018), which are commonly used to manage network data in modeling tasks, may introduce additional perturbations due to their inherent randomness. These challenges underscore the need for a general predictive modeling strategy that can account for network perturbations while ensuring valid inference. Addressing this issue is the central focus of our model development in this paper.
@@ 1.2 Predictive models for network-linked data
We focus on predictive modeling for a node-level response variable. Among existing work, relevant predictive models can be categorized into three main classes based on their design. The first class treats the network as a generalized spatial structure and employs a graph-based autoregressive model to capture dependencies (Zhu et al., 2017; Armillotta and Fokianos, 2023; Wu and Leng, 2023; Chang and Paul, 2024). The second class of methods incorporates network information through a distance-based dependence structure, assuming that responses are independent if the distance between them exceeds a certain threshold (Su et al., 2019; Sit and Ying, 2021; Mukherjee et al., 2021). Both of these classes rely on parametric forms of network effects and assume that the observed networks are accurate. While these methods provide informative model inference if the assumed network effect is appropriate, they may lead to misleading conclusions when the assumed parametric form is violated. Additionally, they tend to be vulnerable to errors in the observed network structure, which is a key issue in our motivating application.
The third class of methods employs nonparametric components to model network effects. For instance, Li et al. (2019) introduces the regression with network cohesion (RNC) approach, which includes an individual node effects component along with a network smoothing penalty. This method has proven to be flexible for modeling network-linked responses and is applicable to various settings, such as generalized linear models. However, this approach lacks a formal statistical inference framework. A more recent model in this category is the subspace linear regression proposed by Le and Li (2022). Instead of assuming a smooth network effect, this model posits that the effect lies within a latent subspace. It offers a valid inference framework and demonstrates robustness against network perturbations. However, the model fitting relies on a sequence of geometric projections, which are valid only for linear regression. This restriction can be limiting in practice, as categorical and discrete responses are common in social science applications, such as our motivating example of the school conflict study. In a separate line of work, Hayes et al. (2022) introduced a model in a similar vein to analyze network-mediated effects in causal problems, in which the network effect is parameterized by a linear combination of latent vectors. This method can handle other types of response variables, but the latent vectors are assumed to follow the random dot product graph model (Athreya et al., 2018).
@@ 2.1 Notations
Throughout the paper, we use $c,C>0$ to denote absolute constants, the values of which may change from line to line. For two sequences of positive scalars $\{a_{n}\}_{n=1}^{\infty}$ and $\{b_{n}\}_{n=1}^{\infty}$, we write $a_{n}=o(b_{n})$ and $a_{n}=O(b_{n})$ if $a_{n}/b_{n}$ converges to zero and $a_{n}/b_{n}$ is bounded, respectively. Similarly, for a sequence of random variables $\{X_{n}\}_{n=1}^{\infty}$, we write $X_{n}=o_{p}(b_{n})$ and $X_{n}=O_{p}(b_{n})$ if $X_{n}/b_{n}$ converges to zero and is bounded in probability, respectively. We use $I_{n}\in\mathbb{R}^{n\times n}$ to denote the identity matrix of size $n$. For a matrix $A=(A_{ij})\in\mathbb{R}^{n\times n}$, $\mathrm{tr}(A)=\sum_{i=1}^{n}A_{ii}$ is the trace, while $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ are the minimum and maximum eigenvalues of $A$, respectively, when $A$ is symmetric. For a vector $u$, $\|u\|$ is the Euclidean norm. For a matrix $W=(W_{ij})\in\mathbb{R}^{m\times n}$ with $1\leq n\leq m$ and the singular value decomposition $W=\sum_{i=1}^{n}\sigma_{i}u_{i}v_{i}^{\top}$, $\|W\|=\max_{1\leq i\leq n}\sigma_{i}$, $\|W\|_{F}=(\sum_{i=1}^{n}\sigma^{2}_{i})^{1/2}$ and $\|W\|_{\infty}=\max_{1\leq i\leq n}\sum_{j=1}^{n}|W_{ij}|$ represent the spectral norm, the Frobenius norm, and the infinity norm of $W$, respectively. In addition, $W_{i}$ is the $i$-th column of $W$, and $W_{u:v}$ is the sub-matrix of $W$ with column vectors $W_{i}$ for $u\leq i\leq v$. We further use $W_{i,u:v}$ to denote the $i$-th row of $W_{u:v}$.
@@ 2.2 Model
We assume there exists a true unobserved relational matrix $P\in\mathbb{R}^{n\times n}$, where $P_{ij}$ describes the strength of the relationship between the nodes $i$ and $j$. Let $\hat{P}=(\hat{P}_{ij})\in\mathbb{R}^{n\times n}$ be an approximate relational matrix, which can be viewed as a noisy version of $P$ that is computable from observed relations between observations. An example from the random‑network modeling literature assumes that the entries of $\hat{P}$ are the observed adjacency connections between nodes, generated as independent Bernoulli random variables with $P=\mathbb{E}[\hat{P}$], or some improved estimators based on certain statistical estimation methods (Li and Le, 2023); another example discussed in detail in Section 4.2 involves stochastic embedding algorithms for which $\hat{P}$ is the similarity between the random embedding output. Intuitively, we expect that $\hat{P}$ does not significantly deviate from $P$.
@@
Conditioning on $X$ and $P$, we assume that $y_{1},...,y_{n}$ are independent random variables drawn from a generalized linear model (GLM). Following McCullagh (2019), the probability density or probability mass function of $y_{i}$ can be expressed in the following form:
@@
Here, $a$, $b$, and $d$ are specific functions depending on the distribution of $y_{i}$. For example, when $a(\phi)=1$, $b(\psi_{i})=\log(1+e^{\psi_{i}})$, and $d(y,\phi)=0$, (1) leads to a logistic regression; when $a(\phi)=1$, $b(\psi_{i})=e^{\psi_{i}}$, and $d(y,\phi)=-\log(y!)$, a Poisson regression is obtained. In addition, $\phi$ is a known dispersion parameter, and $\psi_{i}$ is the natural parameter. We write
@@
through a link function, where $S_{K}(P)$ is the linear subspace spanned by the $K$ leading eigenvectors of $P$. The assumption that $\omega$ belongs to $S_{K}(P)$ is natural and supported by existing evidence that leading eigenvectors of the relational matrix typically capture crucial network information (Özgür et al., 2008; Zeng et al., 2018; Van den Bos et al., 2018; Lee, 2019). In particular, building on the modeling approach outlined in Le and Li (2022), we assume that $\mu$ is contained in the Minkowski sum of $\operatorname{col}(X)$ and $S_{K}(P)$ through the link function $h^{-1}$, which is assumed to be smooth and increasing:
@@ 2.3 Model fitting by the subspace-constrained maximum likelihood
We now describe the model fitting procedure for the network subspace generalized linear model. For ease of presentation, let us first outline this procedure, assuming we observe $S_{K}(P)$. At a high level, we want to use the restricted maximum likelihood estimator (MLE) under the subspace constraint under Definition 2.1. Therefore, the estimation is done by solving the following optimization problem:
@@
$\sigma_{1}=\sigma_{2}=\cdots=\sigma_{r}=1>\sigma_{r+1}\geq\cdots\geq\sigma_{r+s}>0=\sigma_{r+s+1}=\cdots=0,$ (6)
where $r$ and $s$ denote the number of singular values equal to 11 and those taking values strictly between 0 and 1, respectively. Note that it is possible for $r$ and $s$ to be zero. To calculate a basis for the intersection subspace $\mathcal{R}$, let us denote
@@
Viewing $g_{i}$ as a new covariate vector for the $i$-th observation turns the model of Definition 2.1 into a typical generalized linear model with parameter $\gamma^{*}$ (if we do know $g_{i}$’s). Using the first-order stationary condition and setting the gradient of the likelihood function to zero leads to the following estimating equation:
@@
where $h^{\prime}(\cdot)$ is the derivative of the inverse link function and $v(g_{i}^{\top}\gamma)$ is the variance of $y_{i}$. Taking the partial derivative of $S(\gamma)$, we obtain the oracle information matrix
@@
[…]
$\bar{Z}^{\top}\breve{W}=\tilde{U}\tilde{\Sigma}\tilde{V}^{\top}$.
@@
We always assume that $r$, the dimension of $\mathcal{R}$, is known. If it is unknown, Le and Li (2022) proposed a criterion to select $r$ and we can use it here. Specifically, let $\hat{d}=\frac{1}{n}\sum_{i,j=1}^{n}\hat{P}_{ij}$. The following rule can be used to select $r$:
$\hat{r}=\max\left\{i:\hat{\sigma}_{i}\geq 1-\frac{4\sqrt{pK\log n}}{\hat{d}}\right\}$
in which $\hat{\sigma}_{i}$’s are the singular values of $\bar{Z}^{\top}\breve{W}$. Under additional assumptions, Le and Li (2022) showed that $\hat{r}$ can recover $r$ with high probability.
With the known $r$, we estimate $\mathcal{R}$, $\mathcal{C}$, and $\mathcal{N}$ by
$\hat{\mathcal{R}}=\operatorname{col}(\tilde{Z}_{1:r}),\quad\hat{\mathcal{C}}=\operatorname{col}(\tilde{Z}_{(r+1):p}),\quad\hat{\mathcal{N}}=\operatorname{col}(\tilde{W}_{(r+1):K}).$ (15)
@@
1 Calculate the orthonormal basis of $\operatorname{col}(X)$ and form matrix $\bar{Z}\in\mathbb{R}^{n\times p}$ in (7); calculate $K$ eigenvectors of $\hat{P}$ and form $\breve{W}\in\mathbb{R}^{n\times K}$.
2Calculate the singular value decomposition $\bar{Z}^{\top}\breve{W}=\tilde{U}\tilde{\Sigma}\tilde{V}^{\top}$, and form $\tilde{Z}=\sqrt{n}\bar{Z}\tilde{U},\tilde{W}=\sqrt{n}\breve{W}\tilde{V}$.
Find the root $\hat{\gamma}$ of the generalized estimating equation $\tilde{S}(\gamma)=0$ using the iteratively reweighted least squares method, and obtain $\hat{\theta},\hat{\beta},\hat{\alpha}$ by replacing $\gamma$ with $\hat{\gamma}$ in (9), (10), and (11), respectively.
Algorithm 1 Subspace-Constrained Maximum Likelihood Estimation Algorithm